# GroupNL: Low-Resource and Robust CNN Design over Cloud and Device

### v1

@@ Abstract
It has become mainstream to deploy Convolutional Neural Network (CNN) models on ubiquitous Internet of Things (IoT) devices with the help of the cloud to provide users with a variety of high-quality services. Most existing methods have two limitations: (i) low robustness in handling corrupted image data collected by IoT devices; and (ii) high consumption of computational and transmission resources. To this end, we propose the Grouped NonLinear transformation generation method (GroupNL), which generates diversified feature maps by utilizing data-agnostic Nonlinear Transformation Functions (NLFs) to improve the robustness of the CNN model. Specifically, partial convolution filters are designated as seed filters in a convolutional layer, and a small set of feature maps, i.e., seed feature maps, are first generated based on vanilla convolution operation. Then, we split seed feature maps into several groups, each with a set of different NLFs, to generate corresponding diverse feature maps with in-place nonlinear processing. Moreover, GroupNL effectively reduces the parameter transmission between multiple nodes during model training by setting the hyperparameters of NLFs to random initialization and not updating them during model training, and reduces the computing resources by using NLFs to generate feature maps instead of most feature maps generated based on sliding windows. Experimental results on CIFAR-10, GTSRB, CIFAR-10-C, Icons50, and ImageNet-1K datasets in NVIDIA RTX GPU platforms show that the proposed GroupNL outperforms other state-of-the-art methods in model robust and training acceleration. Specifically, on the Icons-50 dataset, the accuracy of GroupNL-ResNet-18 achieves approximately 2.86% higher than the vanilla ResNet-18. GroupNL improves training speed by about 53% compared to vanilla CNN when trained on a cluster of 8 NVIDIA RTX 4090 GPUs on the ImageNet-1K dataset.
@@ 1 Introduction
Internet of Things (IoT) Analytics predicts the number of connected IoT devices worldwide will increase from 16.7 billion in 2023 to 29 billion in 2027111https://iot-analytics.com/number-connected-iot-devices/. It is possible to provide convenient services with the help of connected and ubiquitous IoT devices. Additionally, given the excellent performance of convolutional neural networks (CNN) [1, 2, 3, 4] in computer vision applications, it has become mainstream to deploy CNN models on IoT devices to provide people with various convenient and high-quality services, such as image recognition and monitoring services [5, 6, 7, 8, 9].
@@
IoT devices such as smart cameras and monitors usually work outdoors and often need to process data affected by the natural environment (such as fog, rain, snow, etc.). Vanilla CNN models have low robustness when processing such environmentally affected data. In addition, IoT devices have limited computing and storage resources. Using cloud servers to train CNN models and then sending them to IoT devices for deployment and service provision has become mainstream, as shown in Fig. 1. However, a cloud server usually corresponds to thousands of IoT devices, and frequent deployment and updating of vanilla CNN models will occupy a lot of bandwidth resources. In addition, the vanilla CNN generates feature maps through sliding windows, which consumes a lot of floating-point operations (FLOPs). Therefore, to provide high-quality services on IoT devices, it is urgent to design a CNN model that is highly robust, parameter transmission-friendly and has low computing resource requirements.
This paper aims to propose a CNN design method that increases the diversity of feature maps to improve model robustness without increasing model complexity while ensuring parameter transmission friendliness when deploying and updating models for cloud-assisted IoT devices.
@@
To this end, first, this paper proposes the Grouped NonLinear transformation method, GroupNL. In the GroupNL-based CNN model, partial trainable filters are configured within each convolution layer, which is referred to as seed filters. The feature maps produced via seed filters are called seed feature maps. Subsequently, we split the seed feature maps into many small groups. The remaining feature maps can be generated with diverse nonlinear transformation functions (NLFs) configured with different hyperparameters for each group of seed feature maps. The seed feature maps and generated feature maps are concatenated as the output of the convolutional layer.
GroupNL has four advantages: (i) By grouping seed feature maps, GroupNL can generate diversified feature maps using NLFs with different hyperparameters without increasing the complexity of the model, effectively improving the robustness of the model. (ii) The hyperparameters of NLFs are randomly initialized and remain unchanged during model training. They can be saved and reproduced using random numbers so that after the cloud trains the model, only a small amount of learning parameters and random numbers need to be transmitted. Compared with transmitting the complete model, the amount of parameters transmitted to the IoT device can be effectively reduced. In addition, when the GroupNL-based CNN model employing parallel training technologies [10, 11, 12] like data parallelism (DP) and distributed data parallelism (DDP) to accelerate model training, only a small number of gradients in seed filters have to exchange between multi-GPUs for scatter and gather. This process reduces the intermediate data exchange between different GPUs. Compared with scattering and gathering all gradients in the vanilla CNN model, only exchanging a small number of gradients from seed filters in multi-GPUs memory dramatically accelerates the training speed. (iii) GroupNL includes generating feature maps based on sliding windows and nonlinear transformations. Compared with the vanilla CNN that generates feature maps completely based on sliding windows, it can effectively reduce the number of FLOPs required to run the CNN model. (iv) GroupNL is simple and easy to use, and the convolution modules it generates can directly replace the convolution layers in the existing CNN architecture.
@@
Then, for larger CNN models, such as ResNet-101, this paper introduces sparse technology, that is, apply group convolution to generate seed feature maps, which further reduces the number of learnable parameters and FLOPs of the GroupNL-based model while ensuring model performance. This paper also analyzes four types of NLFs, i.e., Sinusoidal, Monomial, Gaussian, and Laplace functions. It verifies on multiple datasets that the feature maps generated based on the Sinusoidal function have the best performance.
@@
Finally, this paper conducts extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, Icons-50, and ImageNet-1K datasets, and experimental results demonstrate that our proposed GroupNL-based CNN models achieve strong performance in robustness and resource-saving. For example, the accuracy of GroupNL ResNet-34 on CIFAR-10-C is about 2.0% higher than that of vanilla ResNet-34, and the accuracy of GroupNL ResNet-18 on Icons-50 is about 2.86% higher than that of vanilla ResNet-18. We also verify the training acceleration in CIFAR-10 and ImageNet-1K datasets on the NVIDIA RTX 2080Ti/RTX 4090 multi-GPUs platforms. Particularly, with the distributed data parallel training in ImageNet-1K, the GroupNL ResNet-101 achieves an appropriately 53% speedup compared to the vanilla ResNet-101 while maintaining accuracy.
@@
  • This paper proposes the GroupNL, which groups seed feature maps and generates the remaining feature maps using the NLF with different hyperparameters for each group. This improves the diversity of generated feature maps and the robustness of the CNN model without increasing its complexity.
  • This paper introduces sparse technology to the ResNet-101 model, further reducing the number of learnable parameters and FLOPs of the model while maintaining model performance. In addition, it verifies that using the Sinusoidal function is the best to generate feature maps for GroupNL.
  • Extensive experiments on the CIFAR-10, GTSRB, CIFAR-10-C, and Icons-50, and ImageNet-1K datasets demonstrate that our proposed GroupNL achieves strong performance in terms of model robustness, number of learnable parameters, FLOPs, and training speedup.
The remainder of the paper is organized as follows. Section 2 reviews the work on CNN and cloud-assisted CNN training and deployment. Section 3 describes the proposed GroupNL method in detail. Section 4 presents our evaluation results, and Section 5 concludes the paper.
@@ 2 Related Work
This section briefly introduces the Convolutional Neural Network (CNN) models and cloud-device CNN models training and deployment, which are inspiring our work.
@@ 2.1 Convolutional Neural Network Models
Because of the excellent feature extraction capabilities of the convolution operation, many popular CNN models based on convolution operations have been proposed. For example, Simonyan et al. [1] proposed the VGG model through stacked convolution operations with multiple model versions in different layers. He et al. [2] proposed the ResNet with residual branching, which solves the problem of degradation caused by deepening the convolutional layers. Ding et al. [4] introduced the re-parameterization technology to speed up the inference of the CNN model. To deploy CNN models in resource-constrained IoT devices, many lightweight CNN models have been proposed, such as MobileNet [13, 14, 15], ShuffleNet [16, 17], and GhostNet [18], etc. Some methods also consider the robustness of the CNN model, such as MonoCNN [5] and SineFM [19] for cloud-device collaborative deployment.
TABLE I: Feature Maps Generated-based CNN Methods Training Evaluation with ResNet-101 in CIFAR-10 for 300 Epochs on NVIDIA RTX 2080Ti GPUs.
Method Full Conv Nonlinear Func Cheap Linear Norm Top-1 Acc (%) Training Time
Vanilla CNN 95.48 9.3 Hours
MonoCNN 94.97 9.2 Hours
GhostNet 95.04 12.7 Hours
SineFM 95.06 15.4 Hours
GroupNL (Ours) 95.51 8.3 Hours
The most similar to our proposed GroupNL method are the GhostNet [18] and SineFM [19], which also generate feature maps via transformation. However, there are two significant differences as shown in Table I:
@@
(i) The GhostNet generates the feature maps with cheap linear transformation. The cheap linear is a variant of depthwise convolution (DWConv), in which the number of output channels is the multiple of the number of input channels, and the groups of convolution are equal to the number of input channels. So in this paper, we still refer to the cheap linear transformation as “DWConv”. Introducing trainable nodes, i.e., DWConv in the computational graph will result in slower training than vanilla CNN models.
(ii) The SineFM mixes the nonlinear and cheap linear transformation for the feature maps generation, and also introduces batch normalization (Norm) for generated feature maps; SineFM has a better performance compared to GhostNet via nonlinear, nevertheless, it introduces trainable Norm and DWConv resulting in slower training.
Compared to the above methods, our proposed GroupNL is designed with grouped nonlinear transformation to ensure the diversity of generated feature maps and better performance. At the same time, the generation in GroupNL only contains the copying and in-placing operators without extra trainable computational nodes, e.g., DWConv and Norm.
2.2 Cloud-Device CNN Training and Deployment
Cloud-Only. The cloud-only architecture is training and deploying the CNN models on the cloud server. The user sends data and service requests to the cloud server, and after the cloud server processes the data, it sends the results to the users [20, 21, 22, 23]. For example, Liu et al. [23] proposed to deploy the CNN model on the cloud server, and the device obtains recognition services by uploading the collected data to the cloud server. They also process data on edge servers [24, 25, 26, 27, 28, 29] to reduce the amount of data uploaded to the cloud. However, the quality of services of cloud-only architecture is fully constrained by the network condition.
Cloud-Device Collaboration. The second architecture splits the CNN model into two parts, partial CNN layers are deployed on the cloud server and other parts on the IoT devices. The cloud and IoT devices are collaborative training and inference of the split CNN model [30, 31, 32]. For example, Kang et al. [30] proposed to divide the CNN into a head running on the device and a tail running on the cloud and decide the split point based on the load of the device and the conditions of the cloud and model. Stefanos et al. [32] proposed a cloud-device collaborative inference method for the cloud and device with compression technologies to reduce the parameters exchange. However, cloud-device collaborative CNN methods have to search the optimal partitioning points labor-intensively for different CNN models.
Cloud-Assisted Training and Device Deployment. The third architecture is that the cloud server first trains the CNN model and then sends the trained CNN model to the IoT device for deployment [5, 6, 7, 33, 34]. We refer to it as the Cloud-Assisted Architecture. For example, Ding et al. [5] proposed to train a CNN model containing a few learnable parameters on the cloud server and then send the trained CNN model to the IoT device for deployment. Lu et al. [34] proposed to train a lightweight vision transformer model on the cloud server, and then send the trained vision transformer model to the IoT devices for deployment. Furthermore, they reduce the number of model parameters transmitted from the cloud to IoT devices by replacing the attention mechanism with multi-scale pooling technology.
3 Design of the Proposed Approach
3.1 Overview
The proposed GroupNL is based on the cloud-assisted IoT device deployment and updating CNN model architecture. In this architecture, the CNN model is first trained in the cloud server, and then the trained CNN model is transferred to the IoT device for deployment and visual service.
Two key factors motivate us to investigate a robust and low-resource CNN design method.
(i) Typically, IoT devices such as smart surveillance and autonomous driving cameras work outdoors, and the CNN models deployed on the devices often process corrupted data affected by the natural environment, such as images collected on rainy, foggy, and snowy days. We refer to the performance of the on-device CNN model on corrupted data as model robustness. Most existing CNN models have low robustness when dealing with corrupted data.
(ii) IoT devices are resource-constrained. To address the high resource consumption of vanilla CNN models, many alternatives to CNN have been proposed. Although alternatives to the vanilla CNN model reduce the number of model parameters and computation, they contain many additional operations (e.g., DWConv, Norm, etc.), which results in no significant acceleration of model training speed.
To this end, we aim to design a novel, robust, and low-resource CNN model design method to improve the quality of service provided to users by cloud-assisted IoT devices in deploying and updating CNNs.
Figure 2: An example of the GroupNL operation with $c_{\mathrm{in}}=8,c_{\mathrm{out}}=16$, and the number of seed features is $c_{\mathrm{seed}}=4$, the groups for splitting is $g=2$. In this case, the GroupNL can introduce $\gamma\cdot g=6$ different NLFs for the feature maps generation. Compared with the vanilla convolution operation needs 16 filters to compute all features, the GroupNL operation only requires 4 filters to compute the seed features and generates the remaining features by splitting, repeating, and diverse nonlinear transformations.
3.2 Preliminaries of Nonlinear Transformation and Feature Maps Generated-based Methods in CNNs
@@
We illustrate our motivation by analyzing the most relevant works, nonlinear transformation method, i.e., MonoCNN [5], and feature maps generated-based methods, namely GhostNet [18] and SineFM [19], respectively. We also describe the operations of these CNN design methods.
@@
Given the input feature maps ${\bm{x}}$ and output feature maps ${\bm{y}}$ of this convolutional layer, where ${\bm{x}}\in\mathbb{R}^{h\times w\times c_{\mathrm{in}}}$ and ${\bm{y}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times w^{\!{}^{\prime}\!}\times c_{\mathrm{out}}}$. The $h$ and $w$ are the spatial dimensions of the input data/feature maps and $h^{{}^{\prime}}$ and $w^{{}^{\prime}}$ are the spatial dimensions of the output feature maps. $c_{\mathrm{in}}$ is the number of input channels and $c_{\mathrm{out}}$ is the number of output channels.
For a detailed description of the nonlinear transformation and feature maps generated-based methods, we first introduce two concepts from these CNNs:
(i) seed filters refer to a small number of convolutional filters in the convolution layer.
@@
(ii) seed feature maps refer to the feature maps generated by the seed filters and input via a convolution operation.
@@
We denote the seed filters are the ${\bm{W}}_{\mathrm{seed}}$, where ${\bm{W}}_{\mathrm{seed}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{seed}}}$ where $k$ is the spatial dimension of the filter assumed to be square. The number of seed filters is $c_{\mathrm{seed}}$, and the number of generated filters is $c_{\mathrm{gen}}$, where $c_{\mathrm{seed}}$+$c_{\mathrm{gen}}$ = $c_{\mathrm{out}}$. We set the trainable computational operations belonging to $\tt{torch.nn.Module}$ in bold in equation, i.e., $\mathrm{\bf{DWConv}}$ and $\mathrm{\bf{Norm}}$. We also denote a reduction ratio $r$ representing the relationship between $c_{\mathrm{seed}}$ and $c_{\mathrm{out}}$:
$\begin{array}[]{l}r={c_{\mathrm{out}}}/{c_{\mathrm{seed}}}\end{array}.$ (1)
@@
MonoCNN [5]. For a convolutional layer, some filters are designated as seed filters, and the remaining filters are generated based on the seed filters and the data-agnostic nonlinear transformation function (NLF). Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$, and a monomial NLF $\Psi_{\mathrm{mono}}(\cdot)$, we can obtain the remaining filters ${\bm{W}}_{\mathrm{gen}}$ of this layer as follows:
@@
where ${\bm{W}}_{\mathrm{gen}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{gen}}}$. Then, concatenating the seed filter ${\bm{W}}_{\mathrm{seed}}$ and the generated filters ${\bm{W}}_{\mathrm{gen}}$ to obtain the filters ${\bm{W}}$ of this layer:
@@
where ${\bm{W}}\in\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{out}}}$. Finally, perform the vanilla convolution operation on the input data/feature maps to obtain the output ${\bm{y}}$ of this layer:
${\bm{y}}$ $={\bm{W}}*{\bm{x}}$ (4)
$=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})]*{\bm{x}}\ ,$
where * is the convolution operation.

### v2

@@ Abstract
Deploying Convolutional Neural Network (CNN) models on ubiquitous Internet of Things (IoT) devices in a cloud-assisted manner to provide users with a variety of high-quality services has become mainstream. Most existing studies speed up model cloud training/on-device inference by reducing the number of convolution (Conv) parameters and floating-point operations (FLOPs). However, they usually employ two or more lightweight operations (e.g., depthwise Conv, $1\times 1$ cheap Conv) to replace a Conv, which can still affect the model’s speedup even with fewer parameters and FLOPs. To this end, we propose the Grouped NonLinear transformation generation method (GroupNL), leveraging data-agnostic, hyperparameters-fixed, and lightweight Nonlinear Transformation Functions (NLFs) to generate diversified feature maps on demand via grouping, thereby reducing resource consumption while improving the robustness of CNNs. First, in a GroupNL Conv layer, a small set of feature maps, i.e., seed feature maps, are generated based on the seed Conv operation. Then, we split seed feature maps into several groups, each with a set of different NLFs, to generate the required number of diversified feature maps with tensor manipulation operators and nonlinear processing in a lightweight manner without additional Conv operations. We further introduce a sparse GroupNL Conv to speed up by reasonably designing the seed Conv groups between the number of input channels and seed feature maps. Experiments conducted on benchmarks and on-device resource measurements demonstrate that the GroupNL Conv is an impressive alternative to Conv layers in baseline models. Specifically, on Icons-50 dataset, the accuracy of GroupNL-ResNet-18 is 2.86% higher than ResNet-18; on ImageNet-C dataset, the accuracy of GroupNL-EfficientNet-ES achieves about 1.1% higher than EfficientNet-ES. In addition, we verified the efficiency of GroupNL-based models in terms of cloud training and on-device inference.
@@ 1 Introduction
Internet of Things (IoT) Analytics predicts that the number of connected IoT devices worldwide will increase from 16.7 billion in 2023 to 29 billion in 2027111https://iot-analytics.com/number-connected-iot-devices/. It is possible to provide convenient services using connected, ubiquitous IoT devices. Due to the limited computing and storage resources of IoT devices, using the cloud server to train Convolutional Neural Netowrks (CNNs) [35, 13, 18, 6, 16, 34, 17, 8, 36, 7, 37] and then sending them to IoT devices for deployment and service provision (such as image recognition and monitoring services [3, 45, 44, 33, 39]) has become mainstream, as shown in Fig. 1.
@@
Given the limited resources of IoT devices and users’ demand for fast service response, many methods have been proposed to reduce the resource consumption of CNN models and speed up model training/inference by reducing the number of convolutional (Conv) parameters and floating-point operations (FLOPs) during feature map generation. For example, GhostNet [11] uses vanilla Conv and lightweight Conv (e.g., depthwise Conv [16]) to generate feature maps. SineFM [27] combines vanilla Conv, nonlinear transformation functions (NLFs) with fixed hyperparameters, and $1\!\times\!1$ cheap Conv [11] to generate feature maps. These methods can reduce the number of model parameters and FLOPs, but the lightweight operations (depthwise Conv and cheap Conv) introduced affect the model’s cloud training/on-device inference speed. In addition, fixing hyperparameters of NLFs prevents model parameter updates from being completely dependent on the training data, regularizing the model to some extent and improving its robustness, while effectively enhancing the robustness of lightweight CNN models to environmental factors is a task of significant value that remains to be explored.
To this end, this paper aims to propose a CNN design method that improves model robustness by increasing feature map diversity, while reducing model complexity and avoiding the introduction of lightweight operations such as depthwise Conv and cheap Conv to achieve better training/inference speed.
@@
First, this paper proposes the Grouped NonLinear transformation method, GroupNL. In the GroupNL CNN, partially trainable filters are configured within each Conv layer, which are referred to as seed filters or seed Conv. The feature maps produced via seed filters are called seed feature maps. Subsequently, we split the seed feature maps into many small groups. The remaining feature maps can be generated using diverse NLFs, each configured with different hyperparameters for each group of seed feature maps. The hyperparameters of NLFs are fixed, ensuring transmission friendliness when cloud-assisted. The seed feature maps and the generated feature maps are concatenated into the Conv layer’s output. The generation process of GroupNL consists entirely of tensor manipulation operators and nonlinear transformation operations. Compared with GhostNet and SineFM, there is no additional cheap Conv and Batch Normalization. Therefore, GroupNL can achieve better model training/inference acceleration effects.
@@
Then, this paper designs the sparse GroupNL Conv by reasonably setting the relationship between the seed feature map ratio and the number of Conv groups, thereby further reducing the GroupNL model’s parameter count and FLOPs while maintaining performance. This paper also analyzes four types of NLFs: Sinusoidal, Monomial, Gaussian, and Laplace. It verifies across multiple datasets that the feature maps generated using the Sinusoidal function achieve the best performance.
@@
Finally, this paper conducts extensive experiments on the CIFAR-10, CIFAR-10-C, Icons-50, ImageNet-1K, and ImageNet-C datasets, and the experimental results demonstrate that our proposed GroupNL CNN models achieve strong robustness and resource-saving performance. For example, the accuracy of GroupNL ResNet-18 on Icons-50 dataset is about 2.86% higher than that of vanilla ResNet-18. On the ImageNet-C dataset, the accuracy of GroupNL-EfficientNet-ES is appropriately 1.1% higher than the vanilla EfficientNet-ES. In particular, with distributed data parallel training on ImageNet-1K, the GroupNL ResNet-101 achieves a 53% speedup over vanilla ResNet-101 while maintaining accuracy. We also deploy GroupNL models on resource-constrained devices (Raspberry Pi 4B), and on-device measurements show that the GroupNL models outperform in terms of inference speed, memory usage, power, and energy consumption.
@@
  • This paper proposes GroupNL, which groups seed feature maps and uses NLFs with fixed hyperparameters to generate diverse feature maps on demand for each Conv layer. This method improves the diversity of generated feature maps and enhances the CNN model’s robustness without increasing complexity.
  • This paper proposes a sparse mode of GroupNL. By reasonably setting the relationship between the seed feature map ratio and the number of Conv groups, it further reduces the number of parameters and FLOPs of the GroupNL CNN model while ensuring performance.
  • Extensive experiments on the CIFAR-10, CIFAR-10-C, and Icons-50, ImageNet-1K, and ImageNet-C datasets show that our proposed GroupNL achieves impressive performance in terms of model robustness, training/inference speedup.
The remainder of the paper is organized as follows. Section 2 reviews the work on cloud-assisted CNN training and deployment, and CNN. Section 3 describes the proposed GroupNL method in detail. Section 4 presents our evaluation results, and Section 5 concludes the paper.
@@ 2 Related Work
This section briefly introduces cloud-device training and deployment and the Convolutional Neural Network (CNN) models that inspire our work.
2.1 Cloud-Device Training and Deployment
Cloud-Only. The cloud-only architecture trains and deploys the CNN models on the cloud server. The user sends data and service requests to the cloud server, and after the cloud server processes the data, it sends the results to the users [12, 9, 19, 25]. For example, Liu et al. [25] proposed to deploy the CNN model on the cloud server, and the device obtains recognition services by uploading the collected data to the cloud server. They also process data on edge servers [14, 4, 1, 41, 5, 40] to reduce the amount of data uploaded to the cloud. However, the quality of services of cloud-only architecture is fully constrained by the network conditions.
Cloud-Device Collaboration. The second architecture splits the CNN model into two parts: partial CNN layers are deployed on the cloud server, and other parts on the IoT devices. The cloud and IoT devices are collaborative training and inference of the split CNN model [20, 23, 22]. For example, Kang et al. [20] proposed to divide the CNN into a head running on the device and a tail running on the cloud, and decide the split point based on the load of the device and the conditions of the cloud and model. Stefanos et al. [22] proposed a cloud-device collaborative inference method for the cloud and device with compression technologies to reduce the parameter exchange. However, cloud-device collaborative methods have to search for the optimal partitioning points labor-intensively for different CNNs.
Cloud-Assisted Training and Device Deployment. The third architecture is that the cloud server first trains the CNN model and then sends the trained CNN model to the IoT device for deployment [3, 45, 44, 38, 26, 43]. We refer to it as the Cloud-Assisted Architecture. For example, Ding et al. [3] proposed to train a CNN model containing a few learnable parameters on the cloud server and then send the trained CNN model to the IoT device for deployment. Lu et al. [26] proposed to train a lightweight vision transformer model on the cloud server, and then send the trained vision transformer model to the IoT devices for deployment. Furthermore, they reduce the number of model parameters transmitted from the cloud to IoT devices by replacing the attention mechanism with multi-scale pooling technology.
@@ 2.2 Convolutional Neural Network Models
Because of the convolution operation’s excellent feature-extraction capabilities, many popular CNN models have been proposed. For example, Simonyan et al. [35] proposed the VGG model through stacked convolution operations with multiple model versions in different layers. He et al. [13] proposed the ResNet with residual branching, which solves the problem of degradation caused by deepening the convolutional layers. Ding et al. [6] introduced the re-parameterization technology to speed up the inference of the CNN model. To deploy CNN models in resource-constrained IoT devices, many lightweight CNN models have been proposed, such as MobileNet [16, 34, 17], EfficientNet [36, 8, 7, 37], ShuffleNet [46, 28], and GhostNet [11], etc. Some methods also consider the transmission cost and robustness of CNNs, such as SineFM [27] and MonoCNN [3] for cloud-device collaborative deployment.
TABLE I: Comparison with the Existing Convolutions with Conv($n$,$m$,$k$,$g$), height is $h$, width is $w$, and $\xi=\mathrm{Gcd}(n,\frac{n}{r})$.
Method Neural Network Modules #Ops #FLOPs
Vanilla Conv($n$,$m$,kk,1) 1 $whmnk^{2}$
Mono Conv($n$,$m$,kk,1) 1 $whmnk^{2}$
[…]
Depthwise Conv($n$,$n$,$k$,$n$) 1 $whn^{2}k^{2}\frac{1}{n}$
[…]
The most similar methods are Ghost-like Conv, i.e., GhostNet [11] and SineFM [27], which uses a seed Conv to get partial feature maps and leverage a transformation to generate the rest.
@@
  • The GhostNet generates the feature maps with cheap Conv Transformation [11]. The “cheap Conv” is a variant of depthwise Conv [16], in which the number of output channels is a multiple of the number of input channels, and the groups of cheap Conv are equal to the common factor.
  • The SineFM mixes nonlinear functions, Batch Normalization (BN), and cheap Conv for its feature maps generation. It performs better than GhostNet via nonlinearities; however, the introduced BN and cheap Conv result in slower computational speed.
These Ghost-like Conv methods cannot achieve better acceleration because they use many Neural Network Modules, e.g., cheap Conv & BN.
As shown in Table I, for a comparison, we propose GroupNL Conv to solve this issue without using cheap Conv and BN, which only uses a seed Conv with some lightweight Tensor Manipulation Operators like $\tt{torch.cat}$, $\tt{torch.split}$, and $\tt{torch.repeat}$, achieving a better accuracy and speed.
3 The Proposed Method: GroupNL
We introduce GroupNL, a low-resource and robust CNN design method for cloud-assisted IoT deployment, where models are trained in the cloud and deployed on resource-constrained devices. The GroupNL method addresses two key challenges: the demand for real-life resource-efficient training/inference, which prior lightweight alternatives have not fully met, and the need for robustness against real-world corruptions like adverse weather.
We describe the proposed GroupNL in detail as follows: (i) preliminaries in Section 3.1, (ii) detailed design of GroupNL in Section 3.2, and (iii) training and inference acceleration analysis of GroupNL in Section 3.3.
3.1 Preliminaries
@@
We illustrate our motivation by analyzing the most relevant works, the nonlinear transformation method MonoCNN [3], and Ghost-like Conv, including GhostNet [11] and SineFM [27], respectively.
@@
Given the input feature maps ${\bm{x}}\!\in\!\mathbb{R}^{h\times w\times c_{\mathrm{in}}}$ and output feature maps ${\bm{y}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times w^{\!{}^{\prime}\!}\times c_{\mathrm{out}}}$ of this Conv layer, where the $h$ and $w$ are the spatial dimensions of input feature maps, and $h^{{}^{\prime}}$ and $w^{{}^{\prime}}$ for output feature maps. $c_{\mathrm{in}}$, $c_{\mathrm{out}}$ are the number of input and output channels, respectively.
@@
For a detailed description, we introduce two concepts: (i) Seed filters refer to a small number of Conv filters, or referred to as seed Conv; (ii) Seed feature maps refer to feature maps generated by seed filters via seed Conv operation.
@@
We denote seed filters as ${\bm{W}}_{\mathrm{seed}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{seed}}}$, where $k$ is the kernel size. The number of seed filters is $c_{\mathrm{seed}}$, and the number of generated filters is $c_{\mathrm{gen}}$, where $c_{\mathrm{seed}}\!+\!c_{\mathrm{gen}}\!=\!c_{\mathrm{out}}$. We denote an integer reduction ratio $r$ representing the relationship between $c_{\mathrm{seed}}$ and $c_{\mathrm{out}}$, i.e., $r={c_{\mathrm{out}}}/{c_{\mathrm{seed}}}$.
We also set the trainable $\tt{torch.nn.Module}$ operations in bold in equation, e.g., $\mathrm{\bf{Conv}}$ and $\mathrm{\bf{BN}}$.
@@
MonoCNN [3]. For a Conv layer, some filters are designated as seed filters, and the remaining filters are generated based on seed filters and one data-agnostic nonlinear transformation function (NLF). Formally, given seed filters and a monomial NLF $\Psi_{\mathrm{mono}}(\cdot)$, the MonoConv obtains the remaining filters ${\bm{W}}_{\mathrm{gen}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{gen}}}$ as follows:
@@
Then, concatenating the seed filter ${\bm{W}}_{\mathrm{seed}}$ and the generated filters ${\bm{W}}_{\mathrm{gen}}$ to obtain the filters ${\bm{W}}\!\in\!\mathbb{R}^{k\times k\times c_{\mathrm{in}}\times c_{\mathrm{out}}}$:
@@
Finally, perform a vanilla Conv operation on input feature maps to obtain the output $\bm{y}$} of MonoConv:
${\bm{y}}={\bm{W}}*{\bm{x}}=\mathrm{Concat}[{\bm{W}}_{\mathrm{seed}};\Psi_{\mathrm{mono}}({\bm{W}}_{\mathrm{seed}})]*{\bm{x}}\ ,$ (3)